OpenStudy (anonymous):

p=80-Q/2 turn the equation around so that Q is a function of P (that is, add q/2 to both sides and subtract p from both, then solve for q).

OpenStudy (whpalmer4):

You've got the instructions, follow them! \[p = 80 - \frac{q}{2}\]Add \(q/2\) to both sides: \[p+\frac{q}{2} = 80 - \frac{q}{2} + \frac{q}2\]\[p+\frac{q}{2} = 80\]You do the rest...

OpenStudy (anonymous):

p+q=80/2?

OpenStudy (anonymous):

@whpalmer4